# Simple homotopy types of even dimensional manifolds

### v2

@@ 1.6. The involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$
In light of the discussion in Section 1.5, detailed information regarding the involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$ is required in order to complete the proofs of parts (b) and (d) in each of Theorems B, C, LABEL:, and D. The purpose of Part III is to explore this involution in detail and to establish the following three theorems which are the key algebraic ingredients behind the proofs of Theorems B, C, LABEL:, and D respectively.
@@
    |$|\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}|=1$ if and only if $m\in\{2,3,5,6,7,10,11,13,14,17,19\}$; and
@@
    |$|\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}|\to\infty$ super-exponentially in $m$.
Theorem 1.7.
Let $m\geq 2$ be a square-free integer. Then
  1. (i)
    |$|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|=1$ if and only if $m\in\{2,3,5,6,7,10,11,13,14,15,17,19,29\}$; and
  2. (ii)
    |$|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|\to\infty$ super-exponentially in $m$.
@@
    |$|\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}/\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|=1$ for infinitely many $m$; and
@@
    $\sup_{n\leq m}|\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}/\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|\to\infty$ exponentially in $m$.
@@
We will now explain the strategy of proof of these three theorems, as well as some of the key ingredients. The first thing to note is that, whilst $\widetilde{K}_{0}(R)$ is difficult to compute for an arbitrary ring, it is often computable in the case where $R=\mathbb{Z}G$ for a finite group $G$ since finitely generated projective $\mathbb{Z}G$-modules $P$ are all locally free, i.e. $\mathbb{Z}_{p}\otimes_{\mathbb{Z}}P$ is a free $\mathbb{Z}_{p}G$ module for all primes $p$, where $\mathbb{Z}_{p}$ denotes the $p$-adic integers. In particular, $\widetilde{K}_{0}(\mathbb{Z}G)\cong C(\mathbb{Z}G)$ where $C(\cdot)$ denotes the locally free class group (see Section 8.1). The general strategy for determining the involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$ is to note that, since $\widetilde{K}_{0}(\mathbb{Z}C_{m})\cong C(\mathbb{Z}C_{m})$, we can fit $\widetilde{K}_{0}(\mathbb{Z}C_{m})$ into a short exact sequence of abelian groups:
$0\to D(\mathbb{Z}C_{m})\to\widetilde{K}_{0}(\mathbb{Z}C_{m})\to\bigoplus_{d\mid m}C(\mathbb{Z}[\zeta_{d}])\to 0$
@@
where $D(\mathbb{Z}C_{m})$ denotes the kernel group of $\mathbb{Z}C_{m}$ (see Section 8.2) and $C(\mathbb{Z}[\zeta_{d}])$ denotes the ideal class group of $\mathbb{Z}[\zeta_{d}]$, where $\zeta_{d}$ denotes a primitive $d$th root of unity. The standard involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$ restricts to $D(\mathbb{Z}C_{m})$ and induces the involution given by conjugation on each $C(\mathbb{Z}[\zeta_{d}])$ (see Section 10.1).
The proofs of Theorems 1.6 and 1.7 are intertwined and can be found in Section 11.1. The approach we will take is to use Lemma 9.3, and its consequence Lemma 11.4, which allow us to obtain information about the orders |$|\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}|$ and |$|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|$ from the orders of the corresponding groups |$|\{x\in A\mid\overline{x}=-x\}|$ and |$|\{x-\overline{x}\mid x\in A\}|$ in the cases where $A=D(\mathbb{Z}C_{m})$ or $C(\mathbb{Z}[\zeta_{d}]))$ for some $d\mid m$.
@@
The key lemma which makes this approach possible is the following, which is a part of Proposition 10.4. Let $h_{m}=|C(\mathbb{Z}[\zeta_{m}])|$ denotes the class number of the $m$th cyclotomic field, and recall that it splits as a product $h_{m}=h_{m}^{+}h_{m}^{-}$ for integers $h_{m}^{+}=|C(\mathbb{Z}[\zeta_{m}+\zeta_{m}^{-1}])|$ and $h_{m}^{-}$ known as the plus and minus parts of the class number. For an integer $m$, we let $\operatorname{odd}(m)$ denote the odd part of $m$, i.e. the unique odd integer $r$ such that $m=2^{k}r$ for some $k$.
Lemma 1.9.
We have that $\operatorname{odd}(h_{m}^{-})\leq|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|$.
@@
It was shown by Horie [Ho89], using results from Iwasawa theory [Fr82], that there exists finitely many $m$ for which $\operatorname{odd}(h_{m}^{-})=1$ and $\operatorname{odd}(h_{m}^{-})\to\infty$ (see Proposition 10.7). Since Lemma 1.9 also gives a lower bound on |$|\{x\in C(\mathbb{Z}[\zeta_{m}])\mid\overline{x}=-x\}|$, this is enough to prove part (ii) of Theorems 1.6 and 1.7 and to reduce the proof of part (i) to checking finitely many cases. These cases are dealt with via a variety of methods such as analysing the group structure on $C(\mathbb{Z}[\zeta_{m}])$ (see the proof of Proposition 11.6) and the studying the involution on $D(\mathbb{Z}C_{m})$ by relating it to maps between units groups (see Sections 10.4 and 11.1).
The proof of Theorem 1.8 can be found in Section 11.2. Our approach will be based on the isomorphism
$\frac{\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}}{\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}}\cong\widehat{H}^{1}(C_{2};\widetilde{K}_{0}(\mathbb{Z}C_{m}))$
where $\widetilde{K}_{0}(\mathbb{Z}C_{m})$ is viewed as a $\mathbb{Z}C_{2}$-module with the $C_{2}$-action given by the involution (see Proposition 9.5). The short exact sequence above induces a 6-periodic exact sequence on Tate cohomology (Proposition 9.7). This, combined with the fact that $\widehat{H}^{n}(C_{2};A)$ for $A$ finite only depends on its $2$-Sylow subgroup $A_{(2)}$ (Proposition 9.9) gives the following, which is a part of Lemma 11.13.
Lemma 1.10.
If $h_{m}$ is odd, then $\widehat{H}^{1}(C_{2};\widetilde{K}_{0}(\mathbb{Z}C_{m}))\cong\widehat{H}^{1}(C_{2};D(\mathbb{Z}C_{m}))$. In particular, if $h_{m}$ and $|D(\mathbb{Z}C_{m})|$ are both odd, then $\widehat{H}^{1}(C_{2};\widetilde{K}_{0}(\mathbb{Z}C_{m}))=0$.
Using Iwasawa theory, Washington [Wa75] showed that $h_{3^{n}}$ is odd for all $n\geq 1$ (see also work of Ichimura–Nakajima [IN12] which extends this to more primes $p\leq 509$). Since $C_{3^{n}}$ is a $3$-group, $|D(\mathbb{Z}C_{3^{n}})|$ is an abelian $3$-group and so is odd [CR87, Theorem 50.18]. By Lemma 1.10, this implies that $|\widehat{H}^{1}(C_{2};\widetilde{K}_{0}(\mathbb{Z}C_{3^{n}}))|=1$ for all $n\geq 1$ which gives Theorem 1.8 (i).
On the other hand, by Weber’s theorem, $h_{2^{n}}$ is odd for all $n\geq 1$ (see Lemma 10.10). Using a result of Kervaire–Murthy [KM77] which relates $D(\mathbb{Z}C_{2^{n}})$ with $D(\mathbb{Z}C_{2^{n+1}})$ (Proposition 10.19), and the 6-periodic exact sequence on Tate cohomology (Lemma 11.13), we show that $|\widehat{H}^{1}(C_{2};D(\mathbb{Z}C_{2^{n}}))|$ is unbounded as $n\to\infty$. By Lemma 1.10, this implies Theorem 1.8 (ii).
@@
The paper will be structured into three parts. In Part I we develop the necessary background on simple homotopy equivalence, Whitehead torsion and $h$-cobordisms. We will then prove Theorem E, which is our main general result. In Part II we study the manifolds $L\times S^{1}$, leading to the proofs of Theorems B, C, LABEL:, and D subject to results about $\widetilde{K}_{0}(\mathbb{Z}C_{m})$. Part III follows on from Section 1.6: we develop the necessary background on integral representation theory and algebraic number theory. We then study the involution of $\widetilde{K}_{0}(\mathbb{Z}C_{m})$, leading to proofs of Theorems 1.6-1.8.
@@
The following will be in place throughout this article, unless otherwise specified. As above, $n\geq 4$ will be an integer and an $n$-manifold will be a compact connected $\operatorname{CAT}$ $n$-manifold where $\operatorname{CAT}\in\{\operatorname{Diff},\operatorname{PL},\operatorname{TOP}\}$. They will be assumed closed except where it is clear from the context that they are not, e.g. thickenings and cobordisms. We will frequently assume 1.1 but will state this as needed. Groups will be assumed to be finitely presented. Rings $R$ will be assumed to have a multiplicative identity, and $R$-modules will be assumed to be finitely generated.
@@ Acknowledgements
We are grateful to Daniel Kasprowski and Arunima Ray for discussions on the $s$-cobordism theorem, to Henri Johnston, Rachel Newton, and Jack Shotton for advice on ideal class groups, and to Scott Schmieding for discussions on Nil groups.
@@
In this part we will establish general results regarding simple homotopy equivalence. This will be the basis for our applications in Part II. In Section 2, we will recall the basic theory of simple homotopy equivalence, culminating in constraints on the Whitehead torsion of homotopy equivalences between manifolds. Section 3 concerns the two methods which we will use for constructing manifolds: via $h$-cobordisms (Section 3.1) and via the surgery exact sequence (Section 3.2). In Section 4 we study the simple homotopy manifolds sets and prove Theorem E.
@@ 2.1. Simple homotopy equivalence
Let $X$ be a CW complex and let $\phi\colon D^{n}\to X$ be a cellular map. Divide the boundary of the closed $(n+1)$-cell $e^{n+1}$ into two $n$-discs, $\partial e^{n+1}\cong D^{n}\cup_{S^{n-1}}D^{n}$, gluing along the first copy of $D^{n}\subseteq\partial e^{n+1}$. Then the inclusion $X\to X\cup_{\phi}e^{n+1}$ is called an elementary expansion. There is a deformation retract $X\cup_{\phi}e^{n+1}\to X$ in the other direction, and this is called an elementary collapse.
@@
A homotopy equivalence $f\colon X\to Y$ between CW complexes $X$, $Y$ is simple if and only if $\tau(f)=0\in\operatorname{Wh}(\pi_{1}(Y))$.
In dimensions $n\neq 4$, every closed $n$-manifold admits an $n$-dimensional CW structure; see Kirby-Siebenmann [Kirby-Siebenmann:1977-1, III.2.2] for $n\geq 5$, [Ra26] for $n=2$ and [Moise] for $n=3$. In dimension 4, it is an open question whether this holds. Certainly every smooth or PL $n$-manifold admits an $n$-dimensional triangulation, and hence a CW structure.
@@
Let $M$ be a topological $n$-manifold. Embed $M$ in high-dimensional Euclidean space. By [Kirby-Siebenmann:1977-1]*III, §4, there is a normal disc bundle $D(M)\to M$ that admits a triangulation and hence a CW structure. The inclusion map $z_{M}\colon M\to D(M)$ of the 00-section is a homotopy equivalence. Let $z_{M}^{-1}$ denote the homotopy inverse of $z_{M}$.
@@
We say that a homotopy equivalence $f\colon M\to N$ between topological manifolds is simple if the composition $z_{N}\circ f\circ z_{M}^{-1}\colon D(M)\to D(N)$ is simple. Kirby-Siebenmann [Kirby-Siebenmann:1977-1]*III, §4 showed that whether or not this composition is simple does not depend on the choice of normal disc bundle nor on the choice of triangulation.
@@
If $M$, $N$ are smooth or PL, we can ask whether $f$ is simple using the canonical class of triangulations of $M$ and $N$, or by forgetting the smooth/PL structures and using the Kirby-Siebenmann method from Definition 2.3.

### v3

@@ 1.5. The involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$
The purpose of Part III is to explore this involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$ in detail and to prove the following three results which are the key algebraic ingredients behind the proofs of Theorems C, D, LABEL:, and E.
@@
    $|\{y\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{y}=-y\}|=1$ if and only if $m\in\{2,3,5,6,7,10,11,13,14,17,19\}$;
@@
    $|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|=1$ if and only if $m\in\{2,3,5,6,7,10,11,13,14,15,17,19,29\}$; and
  3. (iii)
    $|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|\to\infty$ super-exponentially in $m$, and hence we also have that $|\{y\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{y}=-y\}|\to\infty$ super-exponentially in $m$.
@@
    $|\{y\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{y}=-y\}/\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|=1$ for infinitely many $m$; and
@@
    $\sup_{k\leq m}|\{y\in\widetilde{K}_{0}(\mathbb{Z}C_{k})\mid\overline{y}=-y\}/\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{k})\}|\to\infty$ exponentially in $m$.
@@
We now explain the strategy of proof of these three results, as well as some of the key ingredients. Firstly, whilst $\widetilde{K}_{0}(R)$ is difficult to compute for an arbitrary ring, for a finite group $G$ we have that $\widetilde{K}_{0}(\mathbb{Z}G)\cong C(\mathbb{Z}G)$ where $C(\cdot)$ denotes the locally free class group (see Section 8.1). In the case $G=C_{m}$, this allows us to obtain a short exact sequence:
$0\to D(\mathbb{Z}C_{m})\to\widetilde{K}_{0}(\mathbb{Z}C_{m})\to\bigoplus_{d\mid m}C(\mathbb{Z}[\zeta_{d}])\to 0$ (1.1)
@@
where $D(\mathbb{Z}C_{m})$ denotes the kernel group of $\mathbb{Z}C_{m}$ (see Section 8.2) and $C(\mathbb{Z}[\zeta_{d}])$ denotes the ideal class group of $\mathbb{Z}[\zeta_{d}]$, where $\zeta_{d}:=e^{2\pi i/d}\in\mathbb{C}$. The standard involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$ preserves $D(\mathbb{Z}C_{m})$ and induces the involution given by conjugation on each $C(\mathbb{Z}[\zeta_{d}])$ (see Section 10.1).
@@
To prove Theorem 1.7, we make use of Lemmas 9.3 and 11.3 which allow us to obtain information about the orders $|\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}|$ and $|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|$ from the orders of $|\{x\in A\mid\overline{x}=-x\}|$ and $|\{x-\overline{x}\mid x\in A\}|$ for $A=D(\mathbb{Z}C_{m})$ or $C(\mathbb{Z}[\zeta_{d}]))$ for some $d\mid m$. The key lemma which makes this approach possible is the following, which is Proposition 10.5. Here $h_{m}^{-}$ is a factor of the class number $h_{m}=|C(\mathbb{Z}[\zeta_{m}])|$ (see Definition 10.4). For an integer $m$, we let $\operatorname{odd}(m)$ denote the unique odd integer $r$ such that $m=2^{k}r$ for some $k$.
Proposition 1.9.
$\operatorname{odd}(h_{m}^{-})$ divides $|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|$.
@@
It was shown by Horie [Ho89], using results from Iwasawa theory [Fr82], that there exists finitely many $m$ for which $\operatorname{odd}(h_{m}^{-})=1$ and $\operatorname{odd}(h_{m}^{-})\to\infty$ (see Proposition 10.8). Since Proposition 1.9 also gives a lower bound on $|\{x\in C(\mathbb{Z}[\zeta_{m}])\mid\overline{x}=-x\}|$, this is enough to prove part (iii) of Theorem 1.7 and to reduce the proof of parts (i) and (ii) to checking finitely many cases. These cases are dealt with via a variety of methods such as analysing the group structure on $C(\mathbb{Z}[\zeta_{m}])$ (see the proof of Proposition 11.5) and the studying the involution on $D(\mathbb{Z}C_{m})$ by relating it to maps between units groups (see Sections 10.4 and 11.1).
The proof of Theorem 1.8 can similarly be approached by using (1.1) to relate the Tate cohomology group $\widehat{H}^{1}(C_{2};\widetilde{K}_{0}(\mathbb{Z}C_{m}))$ to the parity of class numbers. This can be found in Section 11.2.
@@
The paper is structured into three parts. In Part I we develop the necessary background on simple homotopy equivalence, Whitehead torsion and $h$-cobordisms. We then prove Theorem B, which is our main general result. In Part II we study the manifolds $L\times S^{1}$, leading to the proofs of Theorems C, D, LABEL:, and E subject to results about $\widetilde{K}_{0}(\mathbb{Z}C_{m})$. In Part III we develop the necessary background on integral representation theory and algebraic number theory. We then study the involution of $\widetilde{K}_{0}(\mathbb{Z}C_{m})$ and prove Theorems 1.7 and 1.8.
@@
The following is in place throughout this article, unless otherwise specified. As above, $n\geq 4$ will be an integer and an $n$-manifold will be a compact connected $\operatorname{CAT}$ $n$-manifold where $\operatorname{CAT}\in\{\operatorname{Diff},\operatorname{PL},\operatorname{TOP}\}$. They will be assumed closed except where it is clear from the context that they are not, e.g. thickenings and cobordisms. We frequently assume 1.1 but will state this as needed. CW complexes will be assumed to be connected. Rings will be assumed to be associative and have a multiplicative identity, but not necessarily commutative.
@@ Acknowledgements
We are grateful to Daniel Kasprowski and Arunima Ray for discussions on the $s$-cobordism theorem, to Henri Johnston, Rachel Newton, and Jack Shotton for advice on ideal class groups, and to Scott Schmieding for discussions on Nil groups. We are extremely grateful to the anonymous referees for their careful reading and insightful comments.
@@
In this part we establish general results regarding simple homotopy equivalence. This is the basis for our applications in Part II. In Section 2, we recall the basic theory of simple homotopy equivalence, culminating in constraints on the Whitehead torsion of homotopy equivalences between manifolds. Section 3 concerns the two methods which we use for constructing manifolds: via $h$-cobordisms (Section 3.1) and via the surgery exact sequence (Section 3.2). In Section 4 we study the simple homotopy manifold sets and prove Theorem B.
@@ 2.1. Simple homotopy equivalence
Let $X$ be a connected CW complex and let $\phi\colon D^{n}\to X$ be a cellular map. Divide the boundary of the closed $(n+1)$-cell $D^{n+1}$ into two $n$-discs, $\partial D^{n+1}\cong D^{n}\cup_{S^{n-1}}D^{n}$, and let $D^{n}\to\partial D^{n+1}$ be the inclusion of the first copy of $D^{n}$. Using this inclusion and $\phi$ we form the union $X\cup_{\phi}D^{n+1}$. The inclusion $X\to X\cup_{\phi}D^{n+1}$ is called an elementary expansion. There is a deformation retract $X\cup_{\phi}D^{n+1}\to X$ in the other direction, and this is called an elementary collapse.
@@
A homotopy equivalence $f\colon X\to Y$ between CW complexes $X$, $Y$ is simple if and only if the element $\tau(f)$ of $\operatorname{Wh}(\pi_{1}(Y))$ is zero.
In dimensions $n\neq 4$, every closed $n$-manifold admits an $n$-dimensional CW structure; see Kirby-Siebenmann [Kirby-Siebenmann:1977-1, III.2.2] for $n\geq 5$, [Ra26] for $n=2$ and [Moise] for $n=3$. In dimension 4, it is an open question whether this holds. Every smooth or PL $n$-manifold admits an $n$-dimensional triangulation, and hence a CW structure.
@@
Let $M$ be a topological $n$-manifold. Embed $M$ in high-dimensional Euclidean space. By [Kirby-Siebenmann:1977-1]*III, §4, there is a normal disc bundle $D(M)\to M$ that admits a triangulation and hence a CW structure. The inclusion map $z_{M}\colon M\to D(M)$ of the 0-section is a homotopy equivalence. Let $z_{M}^{-1}$ denote the homotopy inverse of $z_{M}$.
@@
We say that a homotopy equivalence $f\colon M\to N$ between topological manifolds is simple if the composition $z_{N}\circ f\circ z_{M}^{-1}\colon D(M)\to D(N)$ is simple.
Kirby-Siebenmann [Kirby-Siebenmann:1977-1]*III, §4 showed that whether or not the composition $z_{N}\circ f\circ z_{M}^{-1}$ is simple does not depend on the choice of normal disc bundle nor on the choice of triangulation. In particular, this leads to a well-defined equivalence relation on topological manifolds which extends the notion of simple homotopy equivalence on smooth manifolds.
@@
If $M$, NN are smooth or PL, we can ask whether $f$ is simple using the canonical class of triangulations of $M$ and $N$, or by forgetting the smooth/PL structures and using the Kirby-Siebenmann method from Definition 2.3.